OpenStudy (anonymous):

Graph the equation (x+4)^2/9-(y-7)^2/49=1

OpenStudy (anonymous):

does it look like this? \[\frac{1}{9}(x+4)^2-\frac{1}{49}(y-7)^2=1\]

OpenStudy (anonymous):

no \[\frac{ (x+4)^2 }{ 9 } -\frac{ (y-7)^2 }{ 49 }=1\]

OpenStudy (anonymous):

well it means same thing anyway ^^

OpenStudy (anonymous):

Oh awkward

OpenStudy (anonymous):

is that a hyperbola?

OpenStudy (anonymous):

and that's hyperbola equation! :D never done something similar myself though but seems not so hard and here's link again http://www.mathsisfun.com/geometry/hyperbola.html